Confused????
The height of a triangle is 5 cm less than the length of the base. If the area of the triangle is 117 cm^2, find the height and length of the base?
Thank you

Answer by mananth(16946)

let the length of base be x
height = x-5
area = 117 cm^2
Area of triangle = 1/2 * base * height
117= 1/2 * x(x-5)
2*117=x^2-5x
234=x^2-5x
x^2-5x-234=0
x^2-18x+13x-234=0
x(x-18)+13(x-18)=0
(x-18)(x+13)=0
x=18 OR -13
bur x cannot be negative
base = 18 cm
height = 18-5=13cm